OpenStudy (anonymous):

Simplify log6 4x + log6 5x + log6 2x Possible answers: A) log6 11x B) 3 log6 11x C) log6 40x3 D)3 log6 40x3

myininaya (myininaya):

\[\log(4x \cdot 5x \cdot 2x)=\log(40x^3)\]

OpenStudy (anonymous):

Thanks(: